OpenStudy (anonymous):

A sum of money is divided in ratio 3:5:9.calculate the smallest share given that the largest share is : a. $54 b.$108 c.$369 d.$558

OpenStudy (anonymous):

\[\frac{x}{3}=\frac{54}{9}\] solve for \(x\)

OpenStudy (anonymous):

or just take one third of each number

OpenStudy (anonymous):

why we used 3 and 9 ?

OpenStudy (anonymous):

x=54*3/9 =18 and then ?

OpenStudy (whpalmer4):

If you are dividing the money into 3 shares, in proportion 3:5:9, there are 3+5+9 = 17 parts. Divide the whole amount by 17, then multiply by the number of parts in each share. For example, if the amount to be split was $170, each share would be $170/17 = $10. The first person gets 3 shares = 3*$10 = $30. The second person get 5 shares = 5*$10 = $50. The third person gets 9 shares = 9*$10 = $90. If we add the three together, $30 + $50 + $90 = $170, just as it should.